OpenStudy (sasogeek):

A binary operation \(\large * \) is defined on the set \(\large R \) of real numbers by \(\large x * y = \frac{x+y}{1+xy} \) i. Determine whether or not \(\large * \) is associative ii. Find the identity element \(\large e \)

OpenStudy (anonymous):

What is associative?

OpenStudy (sasogeek):

a*(b*c)=(a*b)*c

OpenStudy (anonymous):

yeah so check it :)

OpenStudy (sasogeek):

well i'm not putting it here to solve it myself lol, i want others to solve it so i see if i made mistakes...

OpenStudy (anonymous):

it is asso...

OpenStudy (anonymous):

but what is identity thingy?

OpenStudy (anonymous):

Ishaan has corroborated!

OpenStudy (anonymous):

<Element> <composition> <identity_element> = <Element> = <identity_element> <composition> <Element>

OpenStudy (sasogeek):

a*e=a

OpenStudy (anonymous):

so you want something which satisfies x*e=x?

OpenStudy (sasogeek):

yeah

OpenStudy (anonymous):

yes but that's incomplete definition saso.

OpenStudy (anonymous):

Left identity and right identity element are unique

OpenStudy (anonymous):

solve it then x+e=e(1+xe)

OpenStudy (anonymous):

(x+e)=x(1+xe)

OpenStudy (anonymous):

0 is the identity element isn't?

OpenStudy (sasogeek):

ok well that's what i got.... 0 :D

OpenStudy (anonymous):

Time to give me medal saso lol

OpenStudy (anonymous):

Additional exercise find the inverse element.

OpenStudy (sasogeek):

i expected that lol, i just started typing the second question :P

OpenStudy (anonymous):

x*e=1/x?

OpenStudy (anonymous):

(x+e)x = (1+xe) x^2 + ex =1 +xe maybe i did something wrong